OpenStudy (anonymous):

For what values of x is the rational expression below undefined? x-7/ 2x^2 -32 Check all that apply. A. 7 B. 4 C. 2 D. -2 E. -4 F. -7

OpenStudy (anonymous):

\[\frac{x-7}{2x^2-32}\] assuming that's the equation that you have. The first thing you have to understand is , what makes a rational expression undefined? One of the key ways you can tell this is when you a value of x makes the expression undefined, or when it makes it so that you divide by zero. Now, looking at your expression, we can simplify it to make it easier to look at. \[\frac{x-7}{2x^2-32}=\frac{x-7}{2(x^2-16)}\]For what values of x will this expression have a 0 on the denominator? You can also write that as \[x^2-16=0\] Solve that problem and you have your answer

OpenStudy (anonymous):

That helps, thank you!